OpenStudy (anonymous):

Find the numerical value of the expression ln(1/e)

OpenStudy (anonymous):

\[\ln = \ln_e\]

OpenStudy (anonymous):

Natural log has its base as \(e\)..

OpenStudy (anonymous):

And : Logarithm property says : \[\ln_a(a) = 1\]

OpenStudy (anonymous):

Also : there is one more property that: \[\ln(a)^m = m \cdot \ln(a)\]

OpenStudy (anonymous):

Also: \[\frac{1}{e} = e^{-1}\]

OpenStudy (anonymous):

Can you make sincere use of all these things? :)